OpenStudy (kimberly_pr):

Compare the functions shown below: f(x) = 4 sin (2x − π) − 1 h(x) = (x − 2)2 + 4

OpenStudy (kimberly_pr):

g(x) x y −1 6 0 1 1 −2 2 −3 3 −2 4 1 5 6

OpenStudy (kimberly_pr):

Compare ^^ those 3 functions And hey @mathmate long time no see, lol

OpenStudy (kimberly_pr):

Oh and the question is: Which function has the smallest minimum y-value? f(x) g(x) h(x) Both f(x) and g(x) have the same minimum y-value.

OpenStudy (kimberly_pr):

^^ That's the question that goes with those functions, I forgot to put it at the beginning, lol

OpenStudy (mathmate):

That's ok. Do you know how to find the minimum of a function?

OpenStudy (kimberly_pr):

Ummm, not really :/

OpenStudy (mathmate):

Have you done calculus yet?

OpenStudy (kimberly_pr):

I'm in algebra 2, but the last module of this class is like and advance to what Calculus is gonna be like, and that's what I'm on now, but it's hard and confusing, to me at least

OpenStudy (kimberly_pr):

*an

OpenStudy (kimberly_pr):

For g(x) is the smallest minimum y-value -3?

OpenStudy (mathmate):

No, that's ok. There are many ways to solve the problem. Telling me what you're doing helps to find the appropriate way.

OpenStudy (mathmate):

As a matter of fact, you don't need calculus to find the minimum of each of the functions. Algebra II is all you need to find the minimum of each. You already gave me correctly the minimum of g(x). Which one do you want to start, f(x) or h(x)?

OpenStudy (kimberly_pr):

Hmm, f(x)

OpenStudy (kimberly_pr):

Would I just plug in a 0 where the x is in the function?

OpenStudy (mathmate):

Not really, for sin(x). ok, f(x)=4sin(2x-\(\pi\))-1 Do you know what is the range of sin(x)?

OpenStudy (kimberly_pr):

From -1 to 1 right?

OpenStudy (mathmate):

Exactly. So what can be the minimum of f(x)?

OpenStudy (kimberly_pr):

Oh, so the minimum is -1 then

OpenStudy (mathmate):

If the minimum of sin(x) (or sin(whatever)) is -1, what is the minimum of f(x)=4sin(2x-pi)-1 ?

OpenStudy (mathmate):

taking into account the minimum of sin(whatever) is -1! :)

OpenStudy (kimberly_pr):

Oh, ok, give me a sec

OpenStudy (kimberly_pr):

I got -1.4, but I feel like it might be wrong.

OpenStudy (mathmate):

not quite 1.4, read on. Or, you can apply the transformation, of the parent function sin(x) to f(x)=4sin(2x-pi)-1 which implies vertical stretch factor of 4, a horizontal shrink of factor 2 (does not affect the minimum) and a vertical translation of -1.

OpenStudy (kimberly_pr):

So the answer for function f(x) is 1.4?

OpenStudy (mathmate):

Well, if you put the min. value of sin(2x-pi)=-1, which is its minimum value, and put it into f(x), tell me what you get.

OpenStudy (kimberly_pr):

So I'd basically do 4 times -1 minus 1? Because that gives me -5

OpenStudy (mathmate):

exactly! So you got the minimum of f(x). Are you ready to move on to h(x)? (...mentally review second degree curves, vertex,....) :)

OpenStudy (kimberly_pr):

Oh yayyy, now I get it. But now what do I do for h(x)?

OpenStudy (mathmate):

h(x) = (x − 2)^2 + 4 It is already in the vertex form, it tells you off the cuff where the vertex is. Can you tell me where the vertex is?

OpenStudy (kimberly_pr):

4

OpenStudy (mathmate):

Actually, the vertex is (2,4). I'll take your answer as correct, since we are working with max/min. Now is the vertex a max or a min? and why?

OpenStudy (kimberly_pr):

Sorry, lol I confused the vertex with the intercept

OpenStudy (kimberly_pr):

Umm I think the vertex is minimum...

OpenStudy (mathmate):

Yes, because when you expand (x-2)^2, the coefficient of x^2 is +1. When the leading coefficient (the coefficient of the highest degree) is positive, the vertex is at minimum. So what is the minimum of h(x)?

OpenStudy (kimberly_pr):

So it's 4, because the 4 is the y, not the 2

OpenStudy (mathmate):

exactly. (2,4) means x=2, y=4. So four is the y-coordinate of the vertex, the minimum. Have you made up your mind about the minimum of h(x)? lol

OpenStudy (kimberly_pr):

Yes, now I understand :) And the answer to the question is f(x) because the minimum of f(x) is -5, which is the lowest

OpenStudy (kimberly_pr):

Thanks!!

OpenStudy (mathmate):

Exactly! you're welcome! :)